# Fuzzy Dark Matter Constraints from the Hubble Frontier Fields

### v1

@@ 6 Current State of FDM Constraints
Figure 6: A selection of constraints on FDM particle mass from various methods and analyses ($2\sigma$ unless otherwise noted). (1) UVLF: This work (Section 3), (Winch et al., 2024; Corasaniti et al., 2017; Schive et al., 2016), (2) cumulative UVLF: This work (Section 4), (Ni et al., 2019), (3) Milky Way satellite galaxies: (Schutz, 2020; Nadler et al., 2021; Banik et al., 2021), (4) Lyman-$\alpha$ forest: (Iršič et al., 2017; Armengaud et al., 2017; Kobayashi et al., 2017; Rogers & Peiris, 2021). See Section 6 or relevant publications for corresponding assumptions and additional details.
@@
First, previous bounds from modeling the luminosity-halo mass relation with the UVLF (Schive et al., 2016; Corasaniti et al., 2017; Winch et al., 2024) are at the $m_{22}\gtrsim 1-3$ level. In this work we improve these bounds by a factor of $\sim 5$ by leveraging the updated Bouwens et al. (2022b) UVLF measurements from the gravitationally magnified galaxies in the HFF. This builds upon many years of study and refinement of the lensed UVLF analyses, including efforts from multiple groups out to $z\sim 9$ (Livermore et al., 2017; Atek et al., 2018; Bhatawdekar et al., 2019; Bouwens et al., 2022b).
Second, we show the similar results attained from the cumulative UVLF by Ni et al. (2019) ($m_{22}>5$ at $3\sigma$). Their determination uses a lensed HFF UVLF analysis from Atek et al. (2018) at $z=6$ compared to a cumulative HMF determined from their own FDM simulations. Their simulations predict fewer halos and thus gives a stronger constraint than the Schive et al. (2016) HMF. [^14] Earlier results from Menci et al. (2017) found $m_{22}>8$ at $3\sigma$, however they relied on UVLF values from an earlier version of Livermore et al. (2017), some of which have since been revised downward (Yung et al., 2018; Ni et al., 2019). Schive et al. (2016) derive a bound of $m_{22}>0.9$ from Bouwens et al. (2015) cumulative UVLF data in the field, i.e. they do not consider the lensed faint-end measurements.
@@
While we believe that this compilation of constraints from the literature is representative, it is by no means complete. Given the broad interest in FDM, it is hard to provide a full summary of the range of techniques being considered. Additional techniques, left out of Figure 6 for simplicity, include bounds from the cosmic microwave background and large scale structure (Hložek et al., 2018), black hole superradiance (Stott & Marsh, 2018; Davoudiasl & Denton, 2019), halo core density profiles (Safarzadeh & Spergel, 2020), ultra-faint dwarf kinematics (Dalal & Kravtsov, 2022) and weak gravitational lensing (Dentler et al., 2022) (see Ferreira (2021) for additional examples and discussion).
@@ 7 Future JWST UVLF
JWST forecasts (Jaacks et al., 2019; Shen et al., 2023) predict detection limits in the field of $M_{\rm UV}\sim-15$. Assuming similar lensing magnification factors to those in the HFF of $\sim 100$ (Atek et al., 2018; Bouwens et al., 2022b), we expect measurements near $M_{\rm UV}\sim-10$ for JWST lensing fields. As shown in Figure 5, the ability to better probe the extreme faint-end of the UVLF may be more important than going deeper in redshift (consider the difference between the blue and black curves contrasted with the values at low and high redshifts).
@@ 8 Conclusion
Similarly to FDM, a warm dark matter (WDM) candidate would suppress small-scale structure (via free-streaming, Bode et al. 2001; Benson et al. 2012; Schneider et al. 2013; Viel et al. 2013.) A rough translation of the FDM constraints to equivalent WDM ones can be performed by matching the relevant half-mode suppression wavenumbers (Marsh & Silk, 2013; Hui et al., 2017). For example, using the thermal relic WDM half-mode wavenumber from Viel et al. (2013), the UVLF modeling (cumulative) constraints of $m_{22}\gtrsim 15(5)$ translate to WDM particle masses of $m_{\rm WDM}\gtrsim 3(2)$ keV. This is competitive with previous determinations using lensed HST UVLF (Corasaniti et al., 2017; Rudakovskyi et al., 2021) ($\sim 2$ keV) and early JWST UVLF measurements (Maio & Viel, 2023; Liu et al., 2024) ($\sim 2-3$ keV). In future work it might be interesting to explore other scenarios, such as non-thermal WDM (e.g. sterile neutrinos (Adhikari et al., 2017; Abazajian, 2017; Boyarsky et al., 2019; Dasgupta & Kopp, 2021)) or interacting dark matter (Tulin & Yu, 2018; Ferreira, 2021; Becker et al., 2021; Adhikari et al., 2022; Dutta & Mahapatra, 2024), which also suppress small-scale structure. Simulations exploring the differences between these theories (and the effects of baryons) will become particularly important if such a suppression is indeed found (Stücker et al., 2022; Shen et al., 2023).
@@
Another uncertain issue relates to whether star formation proceeds similarly in CDM and FDM for intermediate mass halos (Corasaniti et al., 2017; Ni et al., 2019; Shen et al., 2023). Even under the assumptions of the CDM paradigm, the precise role of feedback in the smallest galaxies is an open question (Sun & Furlanetto, 2016; Furlanetto & Mirocha, 2022; Sun et al., 2023; Sipple & Lidz, 2024). Moreover, the surprising number of bright galaxies so far detected by JWST (Harikane et al., 2023; Bouwens et al., 2023; Harikane et al., 2024; Donnan et al., 2024) may indicate that better UVLF modeling at $z\gtrsim 11$ is needed (Mason et al., 2023; Bouwens et al., 2023; Mirocha & Furlanetto, 2023; Muñoz et al., 2024).
@@
However, a much smaller FDM particle mass is possible if it is a subdominant fraction of the total dark matter (for instance $\lesssim 15\%$ being a particle of $m\sim 10^{-24}$ (Winch et al., 2024)). A mixed dark matter (MDM) theory with (at least) one FDM and one CDM component has theoretical motivation in string theory (Arvanitaki et al., 2010). Its existence may also help resolve the $S_{8}$ and other tensions between probes of the matter power spectrum (Amon & Efstathiou, 2022; Rogers et al., 2023; Rogers & Poulin, 2023). For these reasons, understanding the consequences of MDM on the mass function and how it might be tested is an ongoing effort (Marsh & Silk, 2013; Vogt et al., 2023; Chen & Soda, 2023; Laguë et al., 2024; Winch et al., 2024).
@@ Acknowledgements
We thank Raghunath Ghara and Keir Rogers for helpful discussions and David J. E. Marsh for useful email correspondence. JS and AL acknowledge support from NASA ATP grant 80NSSC20K0497. JS, AL, and DG acknowledge support from the Charles Kafuman foundation through grant KA2022-129518. GS was supported by a CIERA Postdoctoral Fellowship.

### v2

@@ 6 Current State of FDM Constraints
Figure 7: A selection of constraints on FDM particle mass from various methods and analyses ($2\sigma$ unless otherwise noted). (1) UVLF: this work (Section 3), (Winch et al., 2024; Corasaniti et al., 2017; Schive et al., 2016), (2) cumulative UVLF: this work (Section 4), (Ni et al., 2019), (3) subhalo mass functions inferred from Milky Way satellite galaxy abundances, strong gravitational lensing, and gaps in stellar streams: (Schutz, 2020; Nadler et al., 2021; Banik et al., 2021), (4) Lyman-$\alpha$ forest: (Iršič et al., 2017; Armengaud et al., 2017; Kobayashi et al., 2017; Rogers & Peiris, 2021). See Section 6 or relevant publications for corresponding assumptions and additional details.
@@
First, previous bounds from modeling the luminosity-halo mass relation with the UVLF (Schive et al., 2016; Corasaniti et al., 2017; Winch et al., 2024) are at the $m_{22}\gtrsim 1-3$ level. In this work we improve these bounds by a factor of $\sim 5$ by leveraging the updated Bouwens et al. (2022b) UVLF measurements from the gravitationally magnified galaxies in the HFF. This builds upon many years of study and refinement of the lensed UVLF analyses, including efforts from multiple groups out to $z\sim 9$ (Livermore et al., 2017; Ishigaki et al., 2018; Atek et al., 2018; Bhatawdekar et al., 2019; Bouwens et al., 2022b).
Second, we show the similar results attained from the cumulative UVLF by Ni et al. (2019) ($m_{22}>5$ at $3\sigma$). Their determination uses a lensed HFF UVLF analysis from Atek et al. (2018) at $z=6$ compared to a cumulative HMF determined from their own FDM simulations. Their simulations predict fewer halos and thus gives a stronger constraint than the Schive et al. (2016) HMF. [^14] Earlier results from Menci et al. (2017) found $m_{22}>8$ at $3\sigma$, however they relied on UVLF values from an earlier version of Livermore et al. (2017), some of which have since been revised downward (Yung et al., 2019; Ni et al., 2019). Schive et al. (2016) derive a bound of $m_{22}>0.9$ from Bouwens et al. (2015) cumulative UVLF data in the field, i.e. they do not consider the lensed faint-end measurements.
@@
While we believe that this compilation of constraints from the literature is representative, it is by no means complete. Given the broad interest in FDM, it is hard to provide a full summary of the range of techniques being considered. Additional techniques, left out of Figure 7 for simplicity, include bounds from the cosmic microwave background and large scale structure (Hložek et al., 2018; Rogers et al., 2023), black hole superradiance (Stott & Marsh, 2018; Davoudiasl & Denton, 2019), halo core density profiles (Safarzadeh & Spergel, 2020), ultra-faint dwarf (UFD) kinematics (Dalal & Kravtsov, 2022; Zimmermann et al., 2024), the size and age of the star cluster in the UFD Eridanus II (Marsh & Niemeyer, 2019), weak gravitational lensing (Dentler et al., 2022), and strong gravitational lensing (Laroche et al., 2022; Powell et al., 2023) (see Ferreira (2021) for additional examples and discussion).
@@ 7 Future JWST UVLF
JWST forecasts (Jaacks et al., 2019; Shen et al., 2024) predict detection limits in the field of $M_{\rm UV}\sim-15$. Assuming similar lensing magnification factors to those in the HFF of $\sim 100$ (Atek et al., 2018; Bouwens et al., 2022b), we expect measurements near $M_{\rm UV}\sim-10$ for JWST lensing fields. As shown in Figure 6, the ability to better probe the extreme faint-end of the UVLF may be more important than going deeper in redshift (consider the difference between the blue and black curves contrasted with the values at low and high redshifts).
@@ 8 Conclusion
Similarly to FDM, a warm dark matter (WDM) candidate would suppress small-scale structure (via free-streaming, Bode et al. 2001; Benson et al. 2013; Schneider et al. 2013; Viel et al. 2013.) A rough translation of the FDM constraints to equivalent WDM ones can be performed by matching the relevant half-mode suppression wavenumbers (Marsh & Silk, 2014; Hui et al., 2017). For example, using the thermal relic WDM half-mode wavenumber from Viel et al. (2013), the UVLF modeling (cumulative) constraints of $m_{22}\gtrsim 15(5)$ translate to WDM particle masses of $m_{\rm WDM}\gtrsim 3(2)$ keV. This is competitive with previous determinations using lensed HST UVLF (Corasaniti et al., 2017; Rudakovskyi et al., 2021) ($\sim 2$ keV) and early JWST UVLF measurements (Maio & Viel, 2023; Liu et al., 2024) ($\sim 2-3$ keV). In future work it might be interesting to explore other scenarios, such as non-thermal WDM (e.g. sterile neutrinos (Adhikari et al., 2017; Abazajian, 2017; Boyarsky et al., 2019; Dasgupta & Kopp, 2021)) or interacting dark matter (Tulin & Yu, 2018; Ferreira, 2021; Becker et al., 2021; Adhikari et al., 2022; Dutta & Mahapatra, 2024; Dave & Goswami, 2024), which also suppress small-scale structure. Simulations exploring the differences between these theories (and the effects of baryons) will become particularly important if such a suppression is indeed found (Stücker et al., 2022; Shen et al., 2024).
@@
Another uncertain issue relates to whether star formation proceeds similarly in CDM and FDM for intermediate mass halos (Corasaniti et al., 2017; Ni et al., 2019; Shen et al., 2024). Even under the assumptions of the CDM paradigm, the precise role of feedback in the smallest galaxies is an open question (Sun & Furlanetto, 2016; Furlanetto & Mirocha, 2022; Sun et al., 2023; Sipple & Lidz, 2024). Moreover, the surprising number of bright galaxies so far detected by JWST (Harikane et al., 2023; Bouwens et al., 2023; Harikane et al., 2024; Donnan et al., 2024) may indicate that better UVLF modeling at $z\gtrsim 11$ is needed (Mason et al., 2023; Bouwens et al., 2023; Mirocha & Furlanetto, 2023; Muñoz et al., 2024).
@@
However, a much smaller FDM particle mass is possible if it is a subdominant fraction of the total dark matter (for instance $\lesssim 15\%$ being a particle of $m\sim 10^{-24}$ (Winch et al., 2024)). A mixed dark matter (MDM) theory with (at least) one FDM and one CDM component has theoretical motivation in string theory (Arvanitaki et al., 2010). Its existence may also help resolve the $S_{8}$ and other tensions between probes of the matter power spectrum (Amon & Efstathiou, 2022; Rogers et al., 2023; Rogers & Poulin, 2023). For these reasons, understanding the consequences of MDM on the mass function and how it might be tested is an ongoing effort (Marsh & Silk, 2014; Vogt et al., 2023; Chen & Soda, 2023; Laguë et al., 2024; Winch et al., 2024).
@@ Acknowledgements
We thank Raghunath Ghara and Keir Rogers for helpful discussions and David J. E. Marsh for useful email correspondence. We thank the anonymous referee for a helpful report. JS and AL acknowledge support from NASA ATP grant 80NSSC20K0497. JS, AL, and DG acknowledge support from the Charles Kafuman foundation through grant KA2022-129518. GS was supported by a CIERA Postdoctoral Fellowship.